The closing price of Schnur Sporting Goods Inc. common stock is uniformly distributed between $18 and $28 per share.

What is the probability that the stock price will be:

  1. More than $20? (Round your answer to 4 decimal places.)

  2. Less than or equal to $26? (Round your answer to 4 decimal places.)
Answer #1

Solution :

Given that,

a = 18

b = 28

(a)

P(x > 20) = (28 - 20) / (28 - 18) = 0.8000

(b)

P(x 26) = (26 - 18) / (28 - 18) = 0.8000